OpenStudy (anonymous):

what is 4(5z-1)-5(z+2)=2(z+1)

OpenStudy (anonymous):

z=16/13

OpenStudy (anonymous):

this is what I got 4(5z-1)-5(z+2)=2(z+1)= 2(z+1)., z= 16/13=1 3/13=1.230769230769231

Directrix (directrix):

@sjackson Do you need the work for this problem?